# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 6. The potential $\mathcal{F}$-essential subgroups
Suppose that $E\in\mathcal{E}(\mathcal{F})$ is non-abelian. Let $T_{E}$ be a Hall $p^{\prime}$-subgroup of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))}(\mathrm{Aut}_{S}(E))$ and $T$ be its lift to $\mathrm{Aut}_{\mathcal{F}}(S)$. Suppose that $V$ is essential in $\mathcal{F}$. Then $T$ normalises $V$ and centralises $C_{V}(S)$. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}}(S)$ is $\mathrm{Aut}(S)$-conjugate to a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S)$, it suffices to analyse the elements of $\Sigma$ which act trivially on $C_{\Lambda(q)}(S_{\Lambda}(q))$. Recall that $C_{\Lambda(q)}(S_{\Lambda}(q))=\langle\overline{xy^{p-1}},\overline{y^{p}}\rangle_{\mathbb{K}}$. Calculating, we see that
$\overline{xy^{p-1}}\cdot\left((1,\theta,\left(\begin{matrix}a&0\\ 0&b\end{matrix}\right)\right)=\theta^{-1}a^{-1}b^{1-p}\overline{xy^{p-1}}\hskip 5.69054pt\mbox{ and }\hskip 5.69054pt\overline{y^{p}}\cdot\left((1,\theta,\left(\begin{matrix}a&0\\ 0&b\end{matrix}\right)\right)=\theta^{-1}b^{-p}\overline{y^{p}}.$
Hence, any element which centralises $C_{\Lambda(q)}(S_{\Lambda}(q))$ satisfies $a=b$ and $\theta=a^{-p}$ and applying Lemma 2.2, we see that such an element must centralise $\Lambda(q)$. This implies that $T$ centralises $V$, a contradiction, and we must have ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$, as required.
If $O_{p}(\mathcal{F})=1$ then since $V$ is characteristic in $S$ we must have ()$\mathcal{E}(\mathcal{F})\neq\{V\}$ by [AKO11, Proposition I.4.5]. Suppose that $\mathcal{E}(\mathcal{F})=R^{S}$. Write $Z:=C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))$. Then for $s\in S$, $Z=C_{R^{s}}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R^{s})))\leq Z(S)$ is normalised by $\mathrm{Aut}_{\mathcal{F}}(S)$ and $\mathrm{Aut}_{\mathcal{F}}(P)$ for $P\in\mathcal{E}(\mathcal{F})$, and so
@@
a contradiction. We conclude that ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$, as desired. ∎
@@
Assume $\{V\}\subsetneq\mathcal{E}(\mathcal{F})$. If $S\cong S_{n}(q)$ then Proposition 6.3 shows that the action of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ is irreducible and $O_{p}(\mathcal{F})\in\{1,V\}$ by [AKO11, Proposition I.4.5]. Since $R<Q$ and ()$\mathcal{E}(\mathcal{F})\cap\{Q,R\}$ is non-empty by Proposition 6.1, it follows that $O_{p}(\mathcal{F})\leq Q$ so that $O_{p}(\mathcal{F})=1$. Similarly if $S\cong S_{\Lambda}(q)$ then ()=$\mathcal{E}(\mathcal{F})=\{R^{S},V\}$ and $O_{p}(\mathcal{F})\leq R\cap V=Z(S)$ by [AKO11, Proposition I.4.5]. Since $C_{V}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V)))=1$ by Proposition 6.3 we conclude that $O_{p}(\mathcal{F})=1$, as desired. ∎
@@
Plainly ()$\mathcal{E}(\mathcal{F})\cup\{S\}\subseteq\mathcal{F}^{frc}$, so it remains to prove the converse. Suppose $P\in\mathcal{F}^{frc}$. If $P<S$ then the Alperin–Goldschmidt theorem implies that an $\mathcal{F}$-conjugate of $P$ is contained in an $\mathcal{F}$-essential subgroup. If all such subgroups are abelian then ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ and $P$ is $\mathcal{F}$-conjugate either to $\{V\}$ or to $R$ since $P$ is $\mathcal{F}$-centric, $V^{\mathcal{F}}=\{V\}$ and $R^{\mathcal{F}}=R^{S}$. We may thus assume that $S=S_{n}(q)$, ()$\mathcal{E}(\mathcal{F})\subseteq\{Q^{S}\}\cup\{V\}$ and some $\mathcal{F}$-conjugate $\widehat{P}$ of $P$ is contained in $Q$. In particular note that $n>1$ and $C_{S}(\widehat{P})\leq\widehat{P}$. Since $Q/Z(Q)$ is a natural module for $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(Q))\cong\operatorname{SL}_{2}(q)$, $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(Q))$ acts transitively on subgroups of order $q$ in $Q/Z(Q)$. If $|P|\leqslant q^{2}$ then let $Y$ be such that $Z(Q)\leq\widehat{P}\leq Y\leq Q$ and $|Y|=q^{2}$; if $|P|>q^{2}$ then let $Y$ be such that $Z(Q)\leq Y\leq\widehat{P}$ with $|Y|=q^{2}$. In either case, $Y$ is $\mathcal{F}$-conjugate to $Z_{2}(S)$. If $\widehat{P}\leq Y$, then $P$ is $\mathcal{F}$-conjugate to a subgroup of $Z_{2}(S)<V$, a contradiction since $V$ centralises $Z_{2}(S)$. Thus, $Z_{2}(S)=Q\cap V$ is properly contained in some $\mathcal{F}$-conjugate of $P$, $P$ is non-abelian and the Alperin–Goldschmidt theorem implies that $P\leq Q^{s}$ for some $s\in S$. If $X\leq P$ is any $\mathrm{Aut}_{\mathcal{F}}(Q^{s})$-conjugate of $Z_{2}(S)$ then $Z(P)=C_{X}(P)=Z(S)=Z(Q^{s})$ so $Q^{s}$ centralises the chain $1\trianglelefteq Z(P)\trianglelefteq P$ and we conclude by Lemma 6.4 that $Q^{s}=P$, as desired. ∎
@@ 7. Uniqueness of the saturated fusion systems on $S$
As in Section 6, here we assume that $p$ is odd, $q=p^{m}$ for some $m>1$, and $\mathcal{F}$ is a core-free saturated fusion system on $S\in S_{n}(q),S_{\Lambda}(q)\}$ for some $11\leqslant n\leqslant p-$ 1. By applying Propositions 6.1 and 6.2, we will determine $\mathcal{F}$ up to isomorphism, thereby completing the proofs of Theorem 1.1 and Theorem 1.2. We start by analysing some small cases.
@@
Plainly $S_{1}(q)$ isomorphic to a Sylow $p$-subgroup of $\mathrm{PSL}_{3}(q)$, and the result follows from [Cle07, Theorem 4.5.1]. ∎
@@
If $S=S_{2}(9)$, $V\in\mathcal{E}(\mathcal{F})$ and $O^{3^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong 2^{.}\mathrm{PSL}_{3}(4)$ then $\mathcal{F}$ is either of the two Henke-Shpectorov exotic fusion system.
@@
This can be verified using the MAGMA package [PS21], and is only a moderate extension of [PS21, Theorem 5.8], which also assumes that $O^{p}(\mathcal{F})=\mathcal{F}$. These fusion systems have the property that ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ and were first discovered in unpublished work of Henke-Shpectorov [HS]. One can prove that they are exotic with an arguments analogous to those in [CP10, Lemma 5.9-Lemma 5.12] for the fusion system $\mathcal{F}(2,9,R)$. ∎
For the remainder of this section, we suppose that $n\geqslant 2$ and that $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{(P)SL}_{2}(q)$ whenever $V\in\mathcal{E}(\mathcal{F})$. The method we use to prove uniqueness is broadly the same for all the saturated fusion systems under consideration.
@@
    If $S\cong S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n,q-1)}$.
@@
    If $S\cong S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n+2,q-1)}$.
@@
    If $S\cong S_{\Lambda}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=(q-1)^{2}$.
@@
We intend to apply [AKO11, Theorem I.7.7] which shows that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is determined by lifting $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))$ for $E\in\mathcal{F}^{frc}$ which is equal to $\{S\}\cup\mathcal{E}(\mathcal{F})$ by Corollary 6.13.
@@
Suppose first that $S\cong S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$. Let $k_{Q}$ be the lift of the element of order $q-1$ in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))}(\mathrm{Aut}_{S}(Q))$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ provided by Lemma 6.10. Then $k_{Q}$ acts on $V$ and $k_{Q}\in C_{\mathrm{Aut}(S)}(Z(S))$. Choose $k$ to be an element of order $\frac{q-1}{(n,2)}$ in $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ which is in the same Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ as $k_{Q}|_{V}$, and chosen such that the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $Q=S_{0}Z_{2}(S)$. Note that when $n$ is even, $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{PSL}_{2}(q)$ so $|C_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))}(C_{V}(S))|=q.\frac{(n,q-1)}{(n,2)}$ by Lemma 2.3. Write $k_{V}$ for the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$. Observe that $K_{Q}:=\langle k_{Q}\rangle$ normalises $K_{V}:=\langle k_{V}\rangle$ and so we need only show that $|K_{Q}\cap K_{V}|=\frac{(n,q-1)}{(n,2)}$ to demonstrate that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n,q-1)}$ as in (1). Since elements of $K_{Q}\cap K_{V}$ act trivially on $C_{V}(S)$, we must have $|K_{Q}\cap K_{V}|\leqslant\frac{(n,q-1)}{(n,2)}$. On the other hand, $C_{K_{V}}(Z(S))$ has order exactly $\frac{(n,q-1)}{(n,2)}$ and normalises $Q$. Since $Q=Z_{2}(S)S_{0}\cong S_{1}(q)$ by Corollary 2.9, Proposition 2.5 implies that $C_{\mathrm{Aut}_{\mathcal{F}}(Q)}(Z(S))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ so that $C_{K_{V}}(Z(S))|_{Q}\leq O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ and then $C_{K_{V}}(Z(S))|_{Q}\leq\langle k_{Q}|_{Q}\rangle$. We obtain $\frac{(n,q-1)}{(n,2)}=|C_{K_{V}}(Z(S))|\leqslant|K_{Q}\cap K_{V}|\leqslant\frac{(n,q-1)}{(n,2)}$.
@@
Next assume that $S\cong S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$. We intend to show that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n+2,q-1)}$ as in (2). Let $k_{R}$ be the lift of the element of order $q-1$ in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ so that $k_{R}$ acts on $V$. Let $k$ be an element of order $\frac{q-1}{(n,2)}$ in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ chosen so that the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $R=S_{0}Z(S)$. We may arrange that $k_{R}|_{V}$ normalises $\langle k\rangle$ and write $k_{V}$ for the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$. The groups $K_{R}:=\langle k_{R}\rangle$ and $K_{V}:=\langle k_{V}\rangle$ can be arranged to commute, so it suffices to prove that $|K_{R}\cap K_{V}|=\frac{(n+2,q-1)}{(n,2)}$. If $1\neq t\in K_{V}$ then $t|_{R}\in K_{R}|_{R}$ if and only if $t_{R}$ acts as $\lambda^{-1}$ on $S/V=RV/V\cong R/Z(S)$ and as $\lambda$ on $Z(S)$ for some $\lambda\in\mathbb{K}^{*}$. Now, $K_{V}$ is the lift of an element of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ and since $V\cong V_{n}(q)|_{\mathbb{F}_{p}}$, we may arrange for $t\in K_{V}$ to act as $\mu^{2}$ on $S/V$ and $\mu^{n}$ on $C_{V}(S)$ for some $\mu\in\mathbb{K}^{*}$. Thus $\lambda^{-1}=\mu^{2}$ and $\lambda=\mu^{n}$ so $\mu^{n+2}=1$ which has exactly $(n+2,q-1)$ solutions, potentially including $\mu=-1$. It follows that $|K_{R}\cap K_{V}|=\frac{(n+2,q-1)}{(n,2)}$, as desired.

### v2

@@ 6. The potential $\mathcal{F}$-essential subgroups
Suppose that $E\in\mathcal{E}(\mathcal{F})$ is non-abelian. Let $T_{E}$ be a Hall $p^{\prime}$-subgroup of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))}(\mathrm{Aut}_{S}(E))$ and $T$ be its lift to $\mathrm{Aut}_{\mathcal{F}}(S)$. Suppose that $V$ is essential in $\mathcal{F}$. Then $T$ normalises $V$ and centralises $C_{V}(S)$. Recall the definition of $P_{\Lambda}^{*}(q)$ from Notation 2.5. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}}(S)$ is $\mathrm{Aut}(S)$-conjugate to a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S)$, it suffices to analyse the elements of $\Sigma$ which act trivially on $C_{\Lambda(q)}(S_{\Lambda}(q))$. Recall that $C_{\Lambda(q)}(S_{\Lambda}(q))=\langle\overline{xy^{p-1}},\overline{y^{p}}\rangle_{\mathbb{K}}$. Calculating, we see that
$\overline{xy^{p-1}}\cdot\left(1,\theta,\left(\begin{matrix}a&0\\ 0&b\end{matrix}\right)\right)=\theta^{-1}a^{-1}b^{1-p}\overline{xy^{p-1}}\hskip 5.69054pt\mbox{ and }\hskip 5.69054pt\overline{y^{p}}\cdot\left(1,\theta,\left(\begin{matrix}a&0\\ 0&b\end{matrix}\right)\right)=\theta^{-1}b^{-p}\overline{y^{p}}.$
Hence, any element which centralises $C_{\Lambda(q)}(S_{\Lambda}(q))$ satisfies $a=b$ and $\theta=a^{-p}$ and applying Lemma 2.6, we see that such an element must centralise $\Lambda(q)$. This implies that $T$ centralises $V$, a contradiction, and we must have $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$, as required.
If $O_{p}(\mathcal{F})=1$ then since $V$ is characteristic in $S$ we must have $\mathcal{E}(\mathcal{F})\neq\{V\}$ by [AKO11, Proposition I.4.5]. Suppose that $\mathcal{E}(\mathcal{F})=R^{S}$. Write $Z:=C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))$. Then for $s\in S$, $Z=C_{R^{s}}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R^{s})))\leq Z(S)$ is normalised by $\mathrm{Aut}_{\mathcal{F}}(S)$ and $\mathrm{Aut}_{\mathcal{F}}(P)$ for $P\in\mathcal{E}(\mathcal{F})$, and so
@@
a contradiction. We conclude that $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$, as desired. ∎
@@
Assume $\{V\}\subsetneq\mathcal{E}(\mathcal{F})$. If $S\cong S_{n}(q)$ then Proposition 6.3 shows that the action of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ is irreducible and $O_{p}(\mathcal{F})\in\{1,V\}$ by [AKO11, Proposition I.4.5]. Since $R<Q$ and $\mathcal{E}(\mathcal{F})\cap\{Q,R\}$ is non-empty by Proposition 6.1, it follows that $O_{p}(\mathcal{F})\leq Q$ so that $O_{p}(\mathcal{F})=1$. Similarly if $S\cong S_{\Lambda}(q)$ then $\mathcal{E}(\mathcal{F})=\{R^{S},V\}$ and $O_{p}(\mathcal{F})\leq R\cap V=Z(S)$ by [AKO11, Proposition I.4.5]. Since $C_{V}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V)))=1$ by Proposition 6.3 we conclude that $O_{p}(\mathcal{F})=1$, as desired. ∎
@@
Plainly $\mathcal{E}(\mathcal{F})\cup\{S\}\subseteq\mathcal{F}^{frc}$, so it remains to prove the converse. Suppose $P\in\mathcal{F}^{frc}$. If $P<S$ then the Alperin–Goldschmidt theorem implies that an $\mathcal{F}$-conjugate of $P$ is contained in an $\mathcal{F}$-essential subgroup. If all such subgroups are abelian then $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ and $P$ is $\mathcal{F}$-conjugate either to $\{V\}$ or to $R$ since $P$ is $\mathcal{F}$-centric, $V^{\mathcal{F}}=\{V\}$ and $R^{\mathcal{F}}=R^{S}$. We may thus assume that $S=S_{n}(q)$, $\mathcal{E}(\mathcal{F})\subseteq\{Q^{S}\}\cup\{V\}$ and some $\mathcal{F}$-conjugate $\widehat{P}$ of $P$ is contained in $Q$. In particular note that $n>1$ and $C_{S}(\widehat{P})\leq\widehat{P}$. Since $Q/Z(Q)$ is a natural module for $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(Q))\cong\operatorname{SL}_{2}(q)$, $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(Q))$ acts transitively on subgroups of order $q$ in $Q/Z(Q)$. If $|P|\leqslant q^{2}$ then let $Y$ be such that $Z(Q)\leq\widehat{P}\leq Y\leq Q$ and $|Y|=q^{2}$; if $|P|>q^{2}$ then let $Y$ be such that $Z(Q)\leq Y\leq\widehat{P}$ with $|Y|=q^{2}$. In either case, $Y$ is $\mathcal{F}$-conjugate to $Z_{2}(S)$. If $\widehat{P}\leq Y$, then $P$ is $\mathcal{F}$-conjugate to a subgroup of $Z_{2}(S)<V$, a contradiction since $V$ centralises $Z_{2}(S)$. Thus, $Z_{2}(S)=Q\cap V$ is properly contained in some $\mathcal{F}$-conjugate of $P$, $P$ is non-abelian and the Alperin–Goldschmidt theorem implies that $P\leq Q^{s}$ for some $s\in S$. If $X\leq P$ is any $\mathrm{Aut}_{\mathcal{F}}(Q^{s})$-conjugate of $Z_{2}(S)$ then $Z(P)=C_{X}(P)=Z(S)=Z(Q^{s})$ so $Q^{s}$ centralises the chain $1\trianglelefteq Z(P)\trianglelefteq P$ and we conclude by Lemma 6.4 that $Q^{s}=P$, as desired. ∎
@@ 7. Uniqueness of the saturated fusion systems on $S$
As in Section 6, here we assume that $p$ is odd, $q=p^{m}$ for some $m>1$, and $\mathcal{F}$ is a core-free saturated fusion system on $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ for some $1\leqslant n\leqslant p-1$. By applying Propositions 6.1 and 6.2, we will determine $\mathcal{F}$ up to isomorphism, thereby completing the proofs of Theorem 1.1 and Theorem 1.3. We start by analysing some small cases.
@@
Lemma 2.101 asserts that $S_{1}(q)$ isomorphic to a Sylow $p$-subgroup of $\mathrm{PSL}_{3}(q)$. The result then follows from [CLE07, Theorem 4.5.1]. ∎
@@
If $S=S_{2}(9)$, $V\in\mathcal{E}(\mathcal{F})$ and $O^{3^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong 2^{.}\mathrm{PSL}_{3}(4)$ then $\mathcal{F}$ is one of the two Henke–Shpectorov exotic fusion system.
@@
This can be verified using the MAGMA package [PS21], and is only a moderate extension of [PS21, Theorem 5.8], which also assumes that $O^{p}(\mathcal{F})=\mathcal{F}$. These fusion systems have the property that $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ and were first discovered in unpublished work of Henke–Shpectorov [18]. One can prove that they are exotic with an argument analogous to those in [CP10, Lemma 5.9-Lemma 5.12] for the fusion system $\mathcal{F}(2,9,R)$. ∎
For the remainder of this section, we suppose that $n\geqslant 2$ and that $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{(P)SL}_{2}(q)$ whenever $V\in\mathcal{E}(\mathcal{F})$. The method we use to prove uniqueness is broadly the same for all the saturated fusion systems under consideration. Recall the fusion systems $\mathcal{F}^{*}_{\Lambda}(q)$ from Notation 3.3, the systems $\mathcal{F}^{*}(n,q,R)$ and $\mathcal{F}^{*}(n,q,Q)$ from Notation 4.1 and the systems $\mathcal{F}^{*}(n,q,R)_{P}$ from Notation 4.4.
@@
    If $S\cong S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n,q-1)}$.
@@
    If $S\cong S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n+2,q-1)}$.
@@
    If $S\cong S_{\Lambda}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is $\mathrm{Aut}(S)$-conjugate to $\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)$ and $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=(q-1)^{2}$.
@@
We intend to apply [AKO11, Theorem I.7.7] which shows that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is determined by lifts of elements in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ for $E\in\mathcal{F}^{frc}$. Observe that $\mathcal{F}^{frc}$ is equal to $\{S\}\cup\mathcal{E}(\mathcal{F})$ by Corollary 6.13.
@@
Suppose first that $S\cong S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$. Let $k_{Q}$ be the lift of the element of order q1q-1 in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))}(\mathrm{Aut}_{S}(Q))$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ provided by Lemma 6.10. Then $k_{Q}$ acts on $V$ and $k_{Q}\in C_{\mathrm{Aut}(S)}(Z(S))$. Choose $k$ to be an element of order $\frac{q-1}{(n,2)}$ in $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ which is in the same Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ as $k_{Q}|_{V}$, and chosen such that the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $Q=S_{0}Z_{2}(S)$. Note that when $n$ is even, $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{PSL}_{2}(q)$ so $|C_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))}(C_{V}(S))|=q.\frac{(n,q-1)}{(n,2)}$ by Lemma 2.7. Write $k_{V}$ for the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$. Observe that $K_{Q}:=\langle k_{Q}\rangle$ normalises $K_{V}:=\langle k_{V}\rangle$ and so we need only show that $|K_{Q}\cap K_{V}|=\frac{(n,q-1)}{(n,2)}$ to demonstrate that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n,q-1)}$ as in (1). Since elements of $K_{Q}\cap K_{V}$ act trivially on $C_{V}(S)$, we must have $|K_{Q}\cap K_{V}|\leqslant\frac{(n,q-1)}{(n,2)}$. On the other hand, $C_{K_{V}}(Z(S))$ has order exactly $\frac{(n,q-1)}{(n,2)}$ and normalises $Q$. Since $Q=Z_{2}(S)S_{0}\cong S_{1}(q)$ by Corollary 2.13, Proposition 2.9 implies that $C_{\mathrm{Aut}_{\mathcal{F}}(Q)}(Z(S))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ so that $C_{K_{V}}(Z(S))|_{Q}\leq O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ and then $C_{K_{V}}(Z(S))|_{Q}\leq\langle k_{Q}|_{Q}\rangle$. We obtain $\frac{(n,q-1)}{(n,2)}=|C_{K_{V}}(Z(S))|\leqslant|K_{Q}\cap K_{V}|\leqslant\frac{(n,q-1)}{(n,2)}$.
@@
Next assume that $S\cong S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$. We intend to show that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=\frac{(q-1)^{2}}{(n+2,q-1)}$ as in (2). Let $k_{R}$ be the lift of the element of order q1q-1 in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ so that $k_{R}$ acts on $V$. Let $k$ be an element of order $\frac{q-1}{(n,2)}$ in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))}(\mathrm{Aut}_{S}(V))$ chosen so that the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $R=S_{0}Z(S)$. We may arrange that $k_{R}|_{V}$ normalises $\langle k\rangle$ and write $k_{V}$ for the lift of $k$ to $\mathrm{Aut}_{\mathcal{F}}(S)$. The groups $K_{R}:=\langle k_{R}\rangle$ and $K_{V}:=\langle k_{V}\rangle$ can be arranged to commute, so it suffices to prove that $|K_{R}\cap K_{V}|=\frac{(n+2,q-1)}{(n,2)}$. If $1\neq t\in K_{V}$ then $t|_{R}\in K_{R}|_{R}$ if and only if $t_{R}$ acts as $\lambda^{-1}$ on $S/V=RV/V\cong R/Z(S)$ and as $\lambda$ on $Z(S)$ for some $\lambda\in\mathbb{K}^{*}$. Now, $K_{V}$ is the lift of an element of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ and since $V\cong V_{n}(q)|_{\mathbb{F}_{p}}$, we may arrange for $t\in K_{V}$ to act as $\mu^{2}$ on $S/V$ and $\mu^{n}$ on $C_{V}(S)$ for some $\mu\in\mathbb{K}^{*}$. Thus $\lambda^{-1}=\mu^{2}$ and $\lambda=\mu^{n}$ so $\mu^{n+2}=1$ which has exactly (n+2,q1)(n+2,q-1) solutions, potentially including $\mu=-1$. It follows that $|K_{R}\cap K_{V}|=\frac{(n+2,q-1)}{(n,2)}$, as desired.